Does the data in the table represent a direct or an inverse variation?
Write an equation to model the data.
x 10, 20, 30, 40, 50
y 15, 7.5, 5, 3.75, 3

Answer by ewatrrr(24785)


 
Hi,
x 10, 20, 30, 40, 50
y 15, 7.5, 5, 3.75, 3
y = k/x |Inverse variation
15 = k/10 |Finding k, the constant of the variation
150 = k
y = 150/x
7.5 = 150/20
5 = 150/30
3.75 = 150/40
3 = 150/50